# Enhancing variational quantum algorithms by balancing training on classical and quantum hardware

### v1

@@ IV Results
Figure 6: XY Hamiltonian VQE (18 Qubits): The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G for 18 qubits. For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared. On the left, we show that both Alternate and Alternate+Simultaneous protocol has better success rates than PSR. For error (Column 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black). The Alternate method only performs well for Layer 1 configuration, which is improved by Alternate+Simultaneous, giving lower error than PSR for all layer configurations. Most importantly, on the right QPU calls reduction is plotted for only Alternate and Alternate+Simultaneous since we compare it relative to PSR (which will coincide with 0% QPU calls reduction). The spread in standard deviation is shown in black. We consistently see reduction in QPU calls using our methods, with frequently lower error values and higher success metrics as compared to PSR.
Figure 7: XY Hamiltonian VQE: The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G in each row, for qubit sizes ranging from 6 to 16. For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared. We consistently see reduction in QPU calls using our methods at high qubits, with frequently lower error values and higher success metrics as compared to PSR.
Figure 8: Gradients for XY Hamiltonian VQE: Variance of the first parameter from PSR and $\mathfrak{g}$-sim block are plotted with increasing qubits for both Full-PSR (blue) and our Alternate method (orange) for finding ground state of XY Hamiltonian using VQE. The diagrams also show the plots for varying number of PSR block layers (1,3,6,9 from top to bottom). For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green. The green curve is the same in all of the plots and provided as a reference for our improved gradients. In both $\mathfrak{g}$-sim and PSR block, we notice a slower decay of gradients as compared to HEA50, with the PSR block maintaining slightly higher magnitude than the $\mathfrak{g}$-sim block. Overall the plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even at high qubits in both of the blocks.
Success: Fraction of the trials that are able to reach below a error threshold of $10^{-3}$, which has been chosen ad-hoc. This gives an estimate of how often can the method get close to the true solution. Going forward, we report the relevant statistics only for the successful trials, except for $\mathfrak{g}$-sim as the error for this case in the mentioned examples is always higher than the threshold.
QPU Calls Reduction: Fraction of quantum circuit evaluations, referred to as Quantum Processing Unit (QPU) calls, reduced in current method as compared to standard training with PSR. We use the number of QPU calls required to reach the point of lowest energy in the corresponding trial as our estimate.
IV.1 Ground State Estimation using VQE
The proposed protocols can be efficiently used for VQE when the Hamiltonian spans a poly-DLA as shown in Sec. IV.1.1. However, for more general cases, where the Hamiltonian spans an exponential DLA, we can still use our method as a pre-training procedure. We discuss this in detail for Longitudinal Transverse Field Ising Model (LTFIM) in Sec. IV.1.2.
IV.1.1 Improved VQE for XY Spin Hamiltonian (poly-DLA)
We apply our chosen ansatz and proposed training schemes on VQE, and demonstrate its effectiveness for this task in comparison to PSR and $\mathfrak{g}$-sim . For this, we consider a XY Spin Hamiltonian given by,
$H_{XY}=\sum_{i}^{N-1}\alpha_{i}\hat{X}_{i}\hat{X}_{i+1}+\beta_{i}\hat{Y}_{i}\hat{Y}_{i+1}$ (14)
XY Spin Hamiltonian has a poly-DLA with the dimension given by $dim(\mathfrak{g})=n^{2}-n$. This is necessary to use $\mathfrak{g}$-sim effectively in our algorithm.
While running the $\mathfrak{g}$-sim -only experiments, all of the DLA elements of the XY spin Hamiltonian produce zero expectation value with the zero initial state $|0^{\otimes n}\rangle$, we added a layer of Hadamard gates $H^{\otimes n}$ to facilitate training . In order to maintain consistency we have added the $H^{\otimes n}$ layer before Block G in all of the examples for the Hamiltonian. Hence, via $\mathfrak{g}$-sim we optimize the following cost function,
$C(\phi)=\langle 0|H^{\otimes n}U_{g}^{\dagger}(\phi)H_{XY}U_{g}(\phi)H^{\otimes n}|0\rangle$ (15)
For the Full-PSR (all parameters of Block Q & G optimized using PSR), Alternate and Alternate + Simultaneous columns we optimize using a larger ansatz by adding the Block Q, and obtain the following cost function,
$C(\theta,\phi)=\langle 0|U_{q}^{\dagger}(\theta)H^{\otimes n}U_{g}^{\dagger}(\phi)H_{XY}U_{g}(\phi)H^{\otimes n}U_{q}(\theta)|0\rangle$ (16)
Now we are not restricted to only searching in the same polynomial space but can search a much larger state space, by virtue of the gates in Block Q. For visualization, we plot the success, error and QPU reduction metrics comparing Full-PSR, Alternate and Alternate + Simultaneous in Fig. 6 and 7, while the actual numbers along with comparison with $\mathfrak{g}$-sim is provided in Table 4 and 3 in Appendix A. The error for $\mathfrak{g}$-sim is significantly higher in the chosen examples, as compared to other methods. Hence, we ignore it in the plots, and only mention the values in Table 3. Later in IV.1.2, we also discuss how this method can be used as a starting point for finding ground states in more general Hamiltonians without the restriction of having a poly-DLA.
Although, we are exploring a larger state space, often the ground state might infact lie close to the space searched by $\mathfrak{g}$-sim , rendering a quantum method unnecessary for the task. This infact happens in the odd qubit examples for XY Hamiltonian, where $\mathfrak{g}$-sim is able get very low errors without additional gates. We mention this here, but in the numerics we only focus on even qubit cases to demonstrate where our method can be effective and standard $\mathfrak{g}$-sim fails. Some of the odd qubit examples are demonstrated in Appendix. B.
Experiment Details: We present the results for 4 different configurations for each qubit, with the number of YZ linear layers: 1,3,6 and 9, changing between the configurations. The results are given in Table 2 and 3 of Appendix A. The Full-PSR column refers to training the full ansatz using standard PSR, Alternate refers to training it by the Alternate training scheme till convergence and Alternate + Simultaneous trains the ansatz using Alternate Scheme for 500 iterations followed by Simultaneous training till convergence. In order to keep the comparison fair, we use an Adam optimizer with initial learning rate $0.01$ and keep it consistent across all experiments.
In Fig. 6 and 7, we have plotted the mean and standard deviation (in black) over the successful trials when measuring the QPU calls reduction. For error, we find the data is often asymmetrically spread around the mean, owing to being lower-bounded by 00 (by definition Eq.(13)) and upper bounded by $10^{-3}$ (success threshold). Hence, for this we resort to the median, and show the spread from $25\%$ to $75\%$ quantiles in black instead of mean and standard deviation. This is due to the fact that median is a more robust metric when the data is distributed asymmetrically. This unequal distribution is also explicit from the $25\%$ to $75\%$ quantiles which extend unequally from the median.
Mitigating Barren Plateaus
We also analyze the gradients of our ansatz for the given VQE task with varying number of qubits. In each case, we track the gradient of the first parameter from each of the Block Q (referred as PSR block) and Block G (referred as $\mathfrak{g}$-sim block) and plots its variance during the training (Fig.8). Each datapoint is obtained after evaluating the variance from the full training of 64 independent trials.
For comparison, we also perform the same VQE task using a 50 layer HEA (referred as HEA50), which is known to exhibit BP [81], and plot the variance of the first parameter of this circuit. In other words, while it is possible that our method will also run into the BP issue at a certain qubit count, atleast for the system sizes we have tested so far, the method holds up well against BP issues in previous methods. Secondly, we believe the reason the direct BP condition in [82, 83] is not applicable here because we choose gates from a poly-DLA and a shallow HEA circuit preventing us from forming a 2-design. However, we do not provide proof of this statement, and propose this as only a potential reason.
HELIA shows a clear improvement in terms of slow decay in gradients in Fig 8. The gradients from $\mathfrak{g}$-sim block demonstrate an exponential decrease, but with much less slope as compared to the HEA50 circuit, while the PSR block shows a sub-exponential decay. Overall, this allows for larger qubit models to be trained efficiently without running into vanishingly small gradients.
IV.1.2 Pre-training in LTFIM Hamiltonian (exp-DLA)
The ability to combine $\mathfrak{g}$-sim and PSR efficiently in our protocol depends on the poly-DLA of the $\mathfrak{g}$-sim block as well as the associated measurement operator. This limits the kind of Hamiltonians on which we can perform VQE task. However, we can still use our method as a starting point for general Hamiltonians with exponential DLA as demonstrated in Algorithm 4.
To start the protocol, we need a step of choosing a subset of operators that form a poly-DLA. In poly-DLA Phase, we perform Alternate, followed by Simultaneous training on the reduced Hamiltonian. During the Exponential DLA Phase of training in Alg.4, we train only the PSR block for some iterations before training the full block. This is done with the aim of keeping the overall quantum resources low, as the PSR block has fewer trainable parameters and hence needs less circuit evaluations to obtain gradients during training. Finally, all of the parameters are trained together using standard PSR.
We demonstrate the results of this protocol for longitudinal-transverse field Ising model (LTFIM),
which has an exponential DLA. A poly-DLA can be constructed by dropping the longitudinal terms
$i\mathfrak{g}_{TFIM}=\langle\{iX_{j}X_{j+1},iZ_{j}\}\rangle_{Lie}$ (18)
(a) LTFIM Hamiltonian with 10 qubits using PSR
(b) LTFIM Hamiltonian with 10 qubits using our proposed method
(c) LTFIM Hamiltonian with 12 qubits using PSR
(d) LTFIM Hamiltonian with 12 qubits using proposed method
Figure 9: Demonstration of PSR (9(a) and 9(c)) and our method (9(b) and 9(d)) to run VQE for LTFIM Hamiltonian (with an exponential DLA) in 10 & 12 qubit cases. The purple and red regions show Alternate and Simultaneous training respectively for the restricted Hamiltonian, followed by training the PSR block and full circuit in green and yellow regions respectively for the full LTFIM Hamiltonian. In both 10 & 12 qubit examples with 3 & 6 YZ linear configurations in the orange block, we notice that our training methods in the initial phase allows VQE to reach much closer to the target ground state during the final training, as compared to just using PSR for the full training, for similar number of of total training iterations. For the overall training, we also slightly reduce the QPU calls ($\sim 6-10\%$)
Experiment Details: The PSR block is composed of varying layers of YZ linear ansatz (3 and 6 layers), while the $\mathfrak{g}$-sim block is generated out of gates from $\mathfrak{g}_{TFIM}$. In the shown examples, we run Alternate optimization for 250 iterations and Simultaneous optimization for 100 iterations with the reduced Hamiltonian, followed by 200 iterations of the PSR block and 1000 iterations of full circuit training. Similarly for Full-PSR training, we run 1450 iterations of PSR to match the number of total iterations in both methods. However, these choices are made ad-hoc and can be modified based on other stopping conditions, such as Early Stopping, that can further reduce QPU calls.
Output: $\theta^{*},\phi^{*}$
Initialize: $\vec{\theta}_{0}\sim\mathcal{N}(0,1);i=0$
begin
       poly-DLA Phase
3       Randomly initialize $\vec{\phi}_{0}\sim\mathcal{N}(0,1)$
4       Define the cost function: $C_{poly}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})H_{poly}U_{g}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$
5       Run Alternate optimization given in Alg.2 for $T_{alt}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
6       Run Simultaneous optimization given in Alg.3 for $T_{sim}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
7       Extract the optimized parameters $(\vec{\theta}^{*},\vec{\phi}^{*})$
       Exponential DLA Phase
8       Define the full cost function, $C_{full}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})HU_{g}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$ initialized at $(\vec{\theta}^{*},\vec{\phi}^{*})$
12
Algorithm 4 VQE for General Hamiltonian
We plot the error metric in Eq.(13) for the LTFIM Hamiltonian with 10 & 12 qubits in Fig.9. Importantly, although the initial training phase is based on the reduced Hamiltonian, the plotted energy values correspond to expectation values of the full Hamiltonian. As shown in Fig.9, the algorithm is indeed able to converge to good solutions by using the our protocols in the first phase of the training.
Compared with PSR, our protocol shows lower error value on an average in Fig. IV.1.2 for both 10 & 12 qubits. More importantly, the QPU calls required for the 10-qubit Hamiltonian VQE using PSR is $54.5e+$ 5 and $56.3e+$ 5 for Layer 3 and 5 respectively, while our proposed method requires $54.065e+$ 5 ($\sim 9.6\%$ reduction) and $55.925e+$ 5 ($\sim 6.5\%$ reduction) QPU calls respectively. Similarly, for 12 qubit, PSR uses $56.12e+$ 5 and $58.28e+$ 5 for Layer 3 and 5 respectively while our method requires $55.466e+$ 5 ($\sim 10.7\%$ reduction) and $57.698e+$ 5 ($\sim 7\%$ reduction) respectively. Overall, our method shows lower error while reducing the required quantum resources.
IV.2 Classification of Quantum Phases
Figure 10: Classification Workflow: $\rho_{i}$ represents the initial state obtained from a training dataset, which is processed using $U_{q}(\theta)$ and $U_{g}(\phi)$ as described before. The measurement operator belongs to a chosen DLA, which also generates the gates for $U_{g}(\phi)$
Our training protocols are useful for QML tasks such as classification. To illustrate this we choose the bond-alternating spin-$1/2$ Heisenberg chain [114],
$H=J\sum_{i}\vec{S}_{2i-1}\cdot\vec{S}_{2i}+J^{\prime}\sum_{i}\vec{S}_{2i}\cdot\vec{S}_{2i+1}$ (19)
We pose the task of learning the quantum phase of the ground state as a classification problem for 10 & 12 qubit examples. Since, quantum phase transition only happens in the limit $N\rightarrow\infty$, it is not accurate to refer to these states as different phases. Rather, we just assign two separate labels ±$\pm 1$ to the ground states for $J<J^{\prime}$ and $J>J^{\prime}$.
Since the qubit counts are small enough, the training and test data (100 train + 100 test dataset) are generated by sampling $J,J^{\prime}$ uniformly from $[-1,1]$ numerically diagonalizing the Hamiltonian in Eq.(19) and obtaining the ground state. This ground state is used as an input into the PQC and the ansatz is trained to generate the correct labels after measuring a Pauli string, that is randomly chosen and fixed at the beginning of the experiment.
For designing the PQC, we encounter the choice of gates for PSR and $\mathfrak{g}$-sim blocks. For the PSR block, we stick to the YZ linear ansatz choice with 9 layers. For $\mathfrak{g}$-sim block however, we do not have a unique choice like we had for previous examples. Hence, we explore 3 different DLA choices to construct the gates:
$i\mathfrak{g}_{XY}$ =$=\langle\{iX_{j}X_{j+1},iY_{j}Y_{j+1}\}\rangle_{Lie}$ (20)
$i\mathfrak{g}_{YZ}$ =$=\langle\{iZ_{j}Z_{j+1},iY_{j}Y_{j+1}\}\rangle_{Lie}$ (21)
$i\mathfrak{g}_{ZX}$ =$=\langle\{iX_{j}X_{j+1},iZ_{j}Z_{j+1}\}\rangle_{Lie}$ (22)
For each of these configurations of DLA, we run 10 independent trials and compare test accuracy. The results are plotted in Fig.1.
We observe different peak accuracies for each of the DLA choices with $i\mathfrak{g}_{YZ}$ providing the highest accuracy in $\mathfrak{g}$-sim for 10 qubits ($0.854\pm 0.008$) and $i\mathfrak{g}_{ZX}$ providing the highest accuracy in $\mathfrak{g}$-sim for 12 qubits ($0.864\pm 0.013$) . In each case, the Alternate and Full-PSR method provide a higher peak accuracy on average, although only by $\sim 1-6\%$. However, the Alternate Method always requires much fewer QPU calls to reach peak accuracy as compared to Full-PSR as shown in Table 1.
In other words, almost same accuracy is reached with our Alternate Method but using much lesser QPU calls compared to only the PSR method, and with better accuracy compared to the $\mathfrak{g}$-sim method.
DLA Full PSR g-sim Alternate
XY 0.898 $\pm$ 0.009 0.860 $\pm$ 0.004 0.897 $\pm$ 0.018
YZ 0.911 $\pm$ 0.008 0.853 $\pm$ 0.005 0.907 $\pm$ 0.013
ZX 0.905 $\pm$ 0.007 0.864 $\pm$ 0.013 0.908 $\pm$ 0.007
(a) Test Accuracy across different methods
DLA Full PSR g-sim Alternate
XY 36736 $\pm$ 30481 - 13381 $\pm$ 6143
YZ 32800 $\pm$ 28216 - 7080 $\pm$ 3540
ZX 25256 $\pm$ 17172 - 11894 $\pm$ 9386
(b) QPU Calls across different methods
Table 1: Classification task using $\mathfrak{g}$-sim , Alternate and Full-PSR Training (with 9 layers in PSR block) for 12 qubits. Three different choices of poly-DLA (referred to as $\mathfrak{g}_{XY},\mathfrak{g}_{YZ}$ and $\mathfrak{g}_{ZX}$) are explored for $\mathfrak{g}$-sim , and also to construct the Block G for Alternate and Full-PSR Method. Average peak test accuracy and number of QPU calls required to reach peak accuracy is reported for each of the configuration. For all examples, the Alternate improves over the accuracy of $\mathfrak{g}$-sim , and is similar to the Full-PSR training with dramatically reduced QPU calls
V Conclusion and Outlook
In order to reach quantum advantage, we need to deploy larger and more complex algorithms on quantum computers while ensuring classical intractability. The reason this is critical for variational algorithms is that, since these are algorithms are heuristic in nature, identifying the thresholds for quantum utility or advantage is usually hard. Therefore, the rather seemingly pedestrian approach of deploying larger algorithms in practice and testing them rigorously, may actually turn out to produce valuable insights.
In this work, we provide a step in that direction by an informed ansatz design and combining $\mathfrak{g}$-sim with PSR, thus delegating the gradient estimation task to both classical and quantum hardware carefully. We extensively test our novel protocols on several practical tasks like VQE and Classification for qubit counts ranging from 6 to 18. In our examples, we notice significant reduction in QPU calls (even upto 60% for the case of VQE in 16 qubit XY Hamiltonian with 1 YZ linear layer), especially at higher number of qubits, alongside improvement in error and success metrics for VQE and accuracy for Classification. Importantly, we also notice a slow decay in gradients as we scale up in qubits, providing a potential strategy in mitigating the ominous Barren Plateau phenomenon. These improvements highlight the benefit of our hybrid gradient-estimation protocol in bringing larger and more complex quantum models closer to us eventually the way towards quantum utility.
  • In the numerical examples, we have restricted ourselves to only Pauli gates as the BCH formula takes a convenient expression in this case. Since in general, this step requires calculating an infinite series, efficient protocols for calculating or approximating it will allow for more interesting gate choices in $\mathfrak{g}$-sim .
  • True quantum utility lies at much higher qubit counts and in using actual quantum hardware. We believe our methods will continue to show improvements beyond 20 qubits, and have more noise-robustness owing to fewer quantum calls required. However, a thorough analysis of the scaling behaviors are necessary.
  • Another direction would be to move away from the standard PQC structure to include non-unitary or general Completely Positive Trace Preserving (CPTP) maps such as in [116].
The authors thank quantum team at Fujitsu Research India Pvt.Ltd., especially Naipunnya Raj, Ruchira Bhat and Rajiv Sangle for their discussions and inputs. The authors extend their immense gratitude to Yasuhiro Endo, Hirotaka Oshima and Shintaro Sato as well as the entire Robust Quantum Computing Department at Fujitsu Limited for their strategic and technical support. The authors also thank Masayoshi Hashima for his valuable support and inputs regarding usage of the state of the art Fujitsu Quantum Simulator. The code used in this paper will be made available at a later date. A preliminary version of this paper has been presented at Fujitsu-IISc Quantum Workshop at Bangalore, India.
Appendix A Numerical Experiments on XY Hamiltonian
Here we tabulate the results of the numerical experiments comparing $\mathfrak{g}$-sim , standard PSR, Alternate and Alternate+Simultaneous method of training on the XY Hamiltonian for various qubit counts in Table 2 and 3, as described in Sec. IV.1.1 and visualized in Fig. 6 and 7. We run VQE on XY Hamiltonian for qubits ranging from 6 to 18, for the configurations with 1,3,6 and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G.
Overall, our method consistently shows reduction in error of estimating the ground state (upto an order of magnitude reduction), improved success rates of reaching correct solution (upto 39% increase) and reduction in QPU calls (upto $\sim 60\%$ reduction), hence reducing the overall quantum resources compared to standard PSR training.
@@ Appendix A Numerical Experiments on XY Hamiltonian
Table 2: XY Hamiltonian (Success and QPU Reduction): The success and QPU calls reduction are shown for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G using Full-PSR, Alternate and Alternate + Simultaneous training (referred as Alt+Sim). The QPU call reduction is measured relative to the Full-PSR method where all parameters are trained by standard PSR. The metrics show that our methods are generally better at reaching higher success rates and lower error(in Table 3), while the QPU calls are reduced significantly
Qubits Error (1e-5)
@@
(a) Error for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
@@
(b) Error for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
@@
(c) Error for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
@@
(d) Error for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Table 3: XY Hamiltonian VQE (Error) : The error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G. The error for $\mathfrak{g}$-sim with only XY Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods. Alternate and Alternate + Simultaneous method shows generally lower errors compared to the Full-PSR training
Appendix B Example instances of $\mathfrak{g}$-sim reaching accurate solution without quantum hardware
Although our aim has been to elevate $\mathfrak{g}$-sim to a hybrid scheme with a more expressive circuit, often it might not be necessary for a given task. In fact, the limited search space might be enough to reach a target solution as can be seen from the below examples. We use $\mathfrak{g}$-sim for finding the ground state of the XY Hamiltonian (Eq. 14) for 13 and 17 qubit cases, and the results are shown in Fig.11. Clearly, $\mathfrak{g}$-sim is able to converge to good solutions (Error $\sim 10^{-3}-10^{-5}$) without adding any additional gates.

### v2

@@ Appendix D Numerical Experiments on XY Hamiltonian
Figure 10: XY Hamiltonian VQE (3,6 and 9 YZ linear layer): The success (Row 1), relative error (Row 2) and QPU calls reduction (Row 3) are plotted for configurations with 3 (a), 6 (b) and 9(c) layer of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G for 6 to 18 qubits. For each qubit count and metric, PSR (orange), Alternate (blue) and Alternate+Simultaneous(purple) training methods are compared. On the first row, we show that Alternate and Alternate+Simultaneous protocol generally has better success rates than PSR. For relative error (Row 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black). The Alternate method struggles in these examples, but Alternate+Simultaneous gives lower relative error than PSR for all qubits. Most importantly, on the lowermost row QPU calls reduction is plotted for only Alternate and Alternate+Simultaneous since we compare it relative to PSR (which will coincide with 0% QPU calls reduction). The spread in standard deviation is shown in black. Except for 6 qubit case in 9 layer configuration, we consistently see reduction in QPU calls using our methods, with lower relative error values and higher success metrics as compared to PSR.
We also tabulate the results of the numerical experiments comparing $\mathfrak{g}$-sim , standard PSR, Alternate and Alternate+Simultaneous method of training on the XY Hamiltonian for various qubit counts in Table 3 and 4, as described in Sec. IV.1.1 and visualized in Fig. 5 and 10. We run VQE on XY Hamiltonian for qubits ranging from 6 to 18, for the configurations with 1, 3, 6 and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G.
Overall, our method consistently shows reduction in relative error of estimating the ground state (upto an order of magnitude reduction), improved success rates of reaching correct solution (upto 39% increase) and reduction in QPU calls (upto $\sim 60\%$ reduction), hence reducing the overall quantum resources compared to standard PSR training.
@@
Table 3: XY Hamiltonian (Success and QPU Reduction): The success and QPU calls reduction are shown for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G using Full-PSR, Alternate and Alternate + Simultaneous training (referred as Alt+Sim). The QPU call reduction is measured relative to the Full-PSR method where all parameters are trained by standard PSR. The metrics show that our methods are generally better at reaching higher success rates and lower relative error(in Table 4), while the QPU calls are reduced significantly
Qubits Relative Error (1e-5)
@@
(a) Relative Error for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Relative Error (1e-5)
@@
(b) Relative Error for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Relative Error (1e-5)
@@
(c) Relative Error for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Relative Error (1e-5)
@@
(d) Relative Error for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Table 4: XY Hamiltonian VQE (Relative Error) : The relative error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G. The relative error for $\mathfrak{g}$-sim with only XY Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods. Alternate and Alternate + Simultaneous method shows generally lower relative error compared to the Full-PSR training
Appendix E Barren Plateau Mitigation for VQE on XY Hamiltonian
Here we plot the variance of the gradients for configurations with 3,6 and 9 YZ linear layers in Block Q of HELIA, when used for VQE on XY Hamiltonian in Fig. 11. In each case, we track the gradient of the first parameter from each of the Block Q (referred as $U_{q}$) and Block G (referred as $U_{g}$) and plots its variance during the training (Fig.11). Each datapoint is obtained after evaluating the variance from the full training of 64 independent trials.
(a) 3 YZ linear layers in Block Q
(b) 6 YZ linear layers in Block Q
(c) 9 YZ linear layers in Block Q
Figure 11: Gradients for XY Hamiltonian VQE: Variance of the first parameter of Block Q ($U_{q}$, red) and G ($U_{g}$, blue) are plotted with increasing qubits for both Full-PSR and our Alternate method for finding ground state of XY Hamiltonian using VQE. The diagrams show the plots for 3(a), 6(b) and 9(c) YZ linear layer in Block Q. For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green. In both Block Q and G, we notice a slower decay of gradients as compared to HEA50. Overall the plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even at high qubits in both of the blocks.
For comparison we plot a deep hardware efficient ansatz (labelled HEA50) in green in Fig.11. HELIA shows a clear improvement in terms of slow decay in gradients for all the configurations of Block Q. The gradients for Block Q and G demonstrate a much slower decay compared to the HEA50 circuit. Overall, this allows for larger qubit models to be trained efficiently without running into vanishingly small gradients.
Appendix F $\mathfrak{g}$-Purity of Hardware Efficient Ansatz at various circuit depths
Using results from Ref. [96] and considering $\mathfrak{g}$ to be simple we obtain Eq.(11). The $\mathfrak{g}$-Purity is defined as,
[…]
Based on Eq. (11), we see that in order to have sub-exponential decay of gradients, $\mathcal{P}_{\mathfrak{g}}(\rho_{q}(\theta))$ or $g$-Purity plays an important role. In Fig. 12, we numerically evaluate it for varying circuit depths considering the $\mathfrak{g}_{XY}$ DLA. Both constant and logarithmic depth circuits show sub-exponential decay while linear depth circuit has exponential decay. This indicates that HELIA with logarithmic depth HEA should also be able to avoid BP. However, we keep that direction open for future exploration.
Figure 12: Average $\mathfrak{g}$-Purity of HEA for $\mathfrak{g}_{XY}$ DLA (defined in Eq. (25)) with increasing qubits and varying circuit depths. For constant ($O(1)$) and logarithmic depth in qubits ($O(log(n))$), the decay is less than exponential (shown in blue and purple respectively). For linear depth ($O(n)$ in green), $\mathfrak{g}$-Purity decreases exponentially. Each point in the plot corresponds to an average of 10000 independent parameter samples.
Appendix G Example instances of $\mathfrak{g}$-sim reaching accurate solution without quantum hardware
Although our aim has been to elevate $\mathfrak{g}$-sim to a hybrid scheme with a more expressive circuit, often it might not be necessary for a given task. In fact, the limited search space might be enough to reach a target solution as can be seen from the below examples. We use $\mathfrak{g}$-sim for finding the ground state of the XY Hamiltonian (Eq. 8) for 13 and 17 qubit cases, and the results are shown in Fig.13. Clearly, $\mathfrak{g}$-sim is able to converge to good solutions (Relative Error $\sim 10^{-3}-10^{-5}$) without adding any additional gates.